What inoculum is needed to get a CFU count of 1.4 x 10^8 CFU/ml when the CFU count is 42, and the dilution factor is 10^-7?
Answer #1

CFU per ml = number of colonies / (volume plated × dilution factor)

Volume plated = innoculum = ? = V

1.4 × 10^8 = 42 / (V × 10^-7)

V = 42 / (1.4 × 10^-7 × 10^8)

V = 3 ml

Innoculum needed is 3 ml.

Please rate.